OpenStudy (anonymous):

Please help: Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> 3.0° 6.0° -7.0° 16.0°

OpenStudy (amistre64):

u.v arccos ( ----------- ) sqrt(u.u * v.v)

OpenStudy (anonymous):

doesnt that make arccos of 1?

OpenStudy (anonymous):

since the sqrt will take off the squared? making the top equal to the bottom

OpenStudy (amistre64):

the sqrt doesnt take off the dot products ...

OpenStudy (amistre64):

u = <2, -4> u = <2, -4> ----------- dot: 4 + 16 = 20 v = <3, -8> v = <3, -8> ----------- dot: 9 + 64 = 73 the sqrt(20*73) is not equal to 20*73

OpenStudy (amistre64):

u = <2, -4> v = <3, -8> ------------ dot: 6 + 32 = 40

OpenStudy (amistre64):

well, 6+32 = 38